OpenStudy (anonymous):

Reggie is three less than four times his daughter's age. Reggie is 69; how old is his daughter? Answer 10 18 20 12

OpenStudy (mr.math):

If x is the daughter's age then 4x-3=69, which implies x=18.

OpenStudy (anonymous):

4x-3=69 4x=69+3 x=72 divided 4 x= 18